6.0 g of cobalt(2) carbonate was added to 40 cm^3 of hydrochloric acid, concentration 2.0 mol/dm3.Calculate the maximum yield of cobalt(2) chloride-6-water and show that the cobalt(2) carbonate was in excess?
1
Expert's answer
2012-09-14T09:52:14-0400
n of CoCO3= 6/119=0.05 moles
n of HCl= 40*2/1000=0.08 moles
using
this equation we can see that only 0.04 moles of CoCO3
0.04 0.08
0.04
CoCO3+2HCl=CoCl2+H2O + CO2

m=0.04*130=5.2
g

CoCl2----------CoCl2*6H2O

m of H2O ----> 6*0.04 *18=
4.32

m of CoCl2*6H2O = 9.52